OpenStudy (anonymous):

If P dollars are deposited at an interest rate r and compounded n times, the future value An can be found by the formula An=P(1+r)^n. Find the rate of interest if a principle amount of $5500 grows to $6655 in two years if interest is compounded annually.

OpenStudy (phi):

Can you assign numbers to An, P, and n in your equation An=P(1+r)^n based on the info in the question?

OpenStudy (anonymous):

So, 6655=5500(1+r)^2. Do I foil the multiply by 5500?

OpenStudy (phi):

starting with 6655=5500(1+r)^2 divide both sides by 5500

OpenStudy (anonymous):

1.21

OpenStudy (phi):

\[\frac{6655}{5500}= (1+r)^2 \] now take the square root of both sides

OpenStudy (phi):

\[ 1.21 = (1+r)^2\]

OpenStudy (anonymous):

\[\sqrt{6655\div5500}\]= 1 + r

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

So, .21?

OpenStudy (phi):

sqrt(1.21)= (1+r) 1.1= 1+r right?

OpenStudy (anonymous):

So, r = .1?

OpenStudy (anonymous):

or 10%

OpenStudy (phi):

yes you can check by using it in your formula: 5500(1+0.1)^2

OpenStudy (phi):

if it is the correct r, you should get 6655

OpenStudy (anonymous):

Your great. Wish I could send you money or something. Many thanks from Georgia.

OpenStudy (phi):

Thanks. btw, you're (you are) not your